James is feeding his horse and he needs a mixture of 88% oats and 12% corn. He currently has 24 pounds of a mixture of 58% oats. How many pounds of oats should be added to the mixture?
Answer by robertb(5830)

0.58*24 + x = 0.88(24 + x),
13.92 + x = 21.12 + 0.88x,
0.12x = 7.2,
x = 60 pounds, the amount of oats he had to add.